# AT2023lli: A Tidal Disruption Event with Prominent Optical Early Bump and Delayed Episodic X-ray Emission

### v1

@@ 3.2 X-ray Spectral Fitting Results
To improve the signal-to-noise ratio of the Swift/XRT spectrum, we combined 9 event files (ObsID: 00016100029, 00016100030, 00016100038–00016100040, 00016100043, 00016100047, 00016100049, 00016100051) with a total exposure time of 17.27 ks. The spectrum is extracted using the xselect tool and fitted with the tbabs*zashift*powerlaw model. The absorption column density is fixed at the Galactic value of (HI4PI Collaboration et al., 2016). This fit results in a photon index of (Cstat/dof=22.73/22) using xspec v12.13.1. A soft X-ray spectrum in TDEs is usually believed to originate from disk emission. Alternatively, the spectrum can be equally well fitted with the tbabs*zashift*diskbb model, yielding a temperature of (Cstat/dof=23.26/22).
To utilize statistics, the XMM-Newton EPIC-pn spectrum is binned with a minimum of 15 photons. However, a single multicolor disk or black body model does not provide a satisfactory fit, leading to a large value. Next, we fit the spectrum using an absorbed power-law model tbabs*zashift*powerlaw. The fit converges to an index (/dof=72.15/57), indicating a soft X-ray spectrum. However, this model did not account for the possible thermal emission from the disk. Therefore, we tried another model tbabs*zashift*(diskbb+powerlaw). This model produced a disk temperature of and a photon index of with /dof=64.61/55. Taking into account the potential intrinsic absorption, we tried to add a photoelectric absorbed component with tbabs*zashift*wabs*(diskbb+powerlaw) in the fitting. The parameters derived for this model were a disk temperature of , a photon index of , and a hydrogen density of (/dof=60.57/54). The power-law component could originate from non-thermal emission from the corona above the disk. To determine whether the absorbed component is necessary, we performed the F-test on the results and derived the F value of 3.6 and the P-value of 0.06. Therefore, we need to be careful about whether or not we need to include this absorbed component. The result is depicted in Figure 5. Although both a power-law model and diskbb model can fit the XRT spectrum well, the latter appears to be more plausible for TDEs because the soft X-ray emission in TDEs likely comes from the disk, which can be better characterized by diskbb. For the XMM spectrum, the additive power-law component may suggest the presence of a corona.
@@ 3.3 Host Galaxy Properties
We collected multiband photometry for the host galaxy of AT2023lli from several archives, including Swift/UVOT, Pan-STARRS, 2MASS, and WISE. We used the Python package Code Investigating GALaxy Emission (CIGALE; Boquien et al. 2019) to model the spectral energy distribution (SED) of the host galaxy. CIGALE can fit the SED of a galaxy from FUV to radio and estimate its physical properties through the analysis of likelihood distribution. We assumed a delayed star formation history (SFH) with an optional exponential burst and used the single stellar population of Bruzual & Charlot (2003). We adopted the dust attenuation curve of Calzetti et al. (2000). Additionally, dust emission is modeled using Dale et al. (2014), and AGN emission is calculated with the model of Stalevski et al. (2012, 2016). The fitting result is shown in Figure 6, which is well fitted solely by stellar components with zero contribution from the AGN component.
@@ 4.1.1 Comparison with Bumps in Other TDEs
Early bumps in the UV/optical light curves are not rare in TDEs that have a well-covered rising phase light curve (Wang et al., 2023b). In Figure 7, we compare the bump observed in AT2023lli with previously reported cases for TDE or TDE candidates. ASASSN-19bt, AT2019azh, and AT2020wey all display bumps that occur within 35 days of the main UV/optical peak (Holoien et al., 2019; Faris et al., 2023; Charalampopoulos et al., 2023). Wang et al. (2023b) reported early bumps that last around 10 days and occur within 30 days from the primary peak in ASASSN-18ap, AT2019mha and AT2019qiz. The latter authors discussed several possibilities for the origin of the bumps, including the emission of unbound debris, the stream-stream collision, the vertical shock compression during the first passage, and the shock breakout of the debris collision. In contrast, the partial TDE candidate ASASSN-14ko recently shows noticeable bumps before the main UV/optical peaks and rebrightenings after that. This can be interpreted as a consequence of the collision between the stream debris and the extended disk in the partial TDE scenario (Huang et al., 2023b), which distinguishes it from AT2023lli and other TDEs.
@@ 4.1.2 Self-crossing of Stream Debris during Circularization and Delayed Accretion?
Thanks to the prominent optical early bump, AT2023lli has been extensively observed in multiwavelength including Swift/XRT, making it a unique object with the earliest X-ray observations among all optical TDEs, starting from 40 days (rest-frame) prior to its peak. It is worth noting that ASASSN-19bt, the only TDE known to possess X-ray observations as early as one month before the peak, also exhibits a tentative early bump in its bolometric light curve (Holoien et al. 2019, see also Figure 7). Interestingly, neither TDE is X-ray detectable throughout the entire rising phase, yet AT2023lli shows delayed X-ray emission in the late declining stage.
We propose a two-phase scenario to explain the optical and X-ray behaviors observed in this unique TDE. A chronological scheme is illustrated in Figure 8. A star is initially tidally disrupted into stream debris, which then undergoes a circularization process to form the accretion disk. Apsidal precession causes streams to self-cross, producing UV/optical emission (Piran et al., 2015). This process may explain the early bump observed in AT2023lli. A new disk forms from the stream debris as it loses energy in the stream-stream collision and circularizes, resulting in the main UV/optical outburst. However, the X-rays from the accretion disk are blocked by the reprocessing layer, probably in the form of collision-induced outflow in the early stage (Lu & Bonnerot, 2020) and super-Eddington-accretion-induced outflow around peak (Roth et al., 2016; Metzger & Stone, 2016; Dai et al., 2018).
@@ 4.2 The Origin of the Episodic X-ray Emission
Taking into account the Keplerian rotation of the optically thin region, its period is , where is the semi-major axis and is the gravitational constant. Assuming the time interval of the two initial X-ray emission as the period, we can estimate the location of the optically thin region as
This result is of the same order of magnitude as the black body radius derived from UV/optical SED fitting, indicating that the “X-ray leak” is located near the outer boundary of the obscured materials. The initial appearance of the X-rays lasted only 2 days, the second lasted 11 days, and the third much longer. This leads us to speculate that if the “X-ray leak” is from the same region, the area is expanding, and we can expect to observe longer-lasting X-ray emission in the future. Subsequent observations confirm this conjecture. A similar scenario was previously used to explain the variability of RXJ1301.9+2747 (Middleton & Ingram, 2015).
@@ 5 Conclusions
The advanced modern time-domain surveys are not only discovering TDEs with an increasing speed but also unveiling the diversity of TDEs in their temporal evolution. AT2023lli, a nearby TDE at redshift 0.036, is exactly such a new member that exhibits notable features in its optical and X-ray light curves. In this study, we conducted continuous multiwavelength observations of it. These observations revealed characteristic properties, including an almost constant blackbody temperature, very broad Balmer and possible He II 4686 Å emission. However, it also exhibits some distinctive features that have captured our attention. The UV/optical light curves of AT2023lli show a prominent bump that lasts nearly a month in the early rising phase. On the other hand, the time interval between the bump peak and the primary peak is nearly two months, making both timescales the longest among all TDEs with detected bumps, albeit only a handful so far. X-ray emission was not significantly detected until the late UV/optical decline stage ( days post-peak), despite two weak episodic emissions earlier. The XMM-Newton ToO observation of AT2023lli, triggered by us upon persistent X-ray detection, revealed X-ray spectra with a possible intrinsic absorption component, albeit with large uncertainties. It is worth noting that AT2023lli exhibits the power-law decay with an index of -4.10, making it one of the fastest declining optically-selected TDEs so far.
Taking all these observations into account, we suggest that the UV/optical bump may be caused by the self-crossing of the debris streams of the TDE, while the delayed accretion produced the UV/optical main outburst. The self-crossing or circularization of the debris streams produce outflows, which surround the accretion disk and cover it, leading to the absorption of soft X-ray photons from the inner disk and the detection of only the UV/optical emission. Due to the inhomogeneous distribution of the obscuring material, some regions were optically thin, allowing some X-ray photons to “leak” in the decline phase of the UV/optical. As the obscuring material evolved to be optically thin, the X-ray photons could be gradually detected. We have also discussed the possibility of double TDEs, in which the two stars in the binary are disrupted in sequence by the SMBH, which is also a possible scenario.
@@
We thank the anonymous referee for providing valuable comments, which help to improve the manuscript. This work is supported by National Key Research and Development Program of China (2023YFA1608100), the National Natural Science Foundation of China (grants 12393814, 12073025,12192221,12233008,12025303), the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0550200, XDB41000000, XDB0550300), Anhui Provincial Natural Science Foundation (2308085QA32), the China Manned Space Project (No. CMS-CSST-2021-A13) and the Fundamental Research Funds for Central Universities (WK3440000006). L.L. acknowledges the support of the National Natural Science Foundation of China (grant No. 12303050), the China Postdoctoral Science Foundation (2023M743397), and the Fundamental Research Funds for the Central Universities. The authors appreciate the support of the Cyrus Chun Ying Tang Foundations. The authors express their gratitude to Ping Cai for her help in creating the figure. We also appreciate the members of the WFST operation and maintenance team for their support. The Wide Field Survey Telescope (WFST) is a joint facility of the University of Science and Technology of China, Purple Mountain Observatory. This research uses data obtained through the Telescope Access Program (TAP), which has been funded by the TAP member institutes. Observations with the Hale Telescope at Palomar Observatory were obtained as part of an agreement between the National Astronomical Observatories, the Chinese Academy of Sciences, and the California Institute of Technology. The authors acknowledge the support of the Lijiang 2.4m telescope staff. Funding for the telescope has been provided by the Chinese Academy of Sciences and the People’s Government of Yunnan Province”. The authors acknowledge the use of public data from the Swift data archive. The authors thank the Swift ToO team for accepting our proposal and executing the observations. Based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. This work makes use of observations from the Las Cumbres Observatory global telescope network. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile.

### v2

@@ 3.2 X-Ray Spectral Fitting Results
To improve the S/N ratio of the Swift/XRT spectrum, we combined nine event files (ObsID: 00016100029, 00016100030, 00016100038–00016100040, 00016100043, 00016100047, 00016100049, 00016100051) with a total exposure time of 17.27 ks. The spectrum is extracted using the xselect tool and fitted with the tbabs*zashift*powerlaw model. The absorption column density is fixed at the Galactic value of (HI4PI Collaboration et al., 2016). This fit results in a photon index of (Cstat/dof=22.73/22) using xspec v12.13.1. A soft X-ray spectrum in TDEs is usually believed to originate from disk emission. Alternatively, the spectrum can be equally well fitted with the tbabs*zashift*diskbb model, yielding a temperature of (Cstat/dof=23.26/22).
To utilize statistics, the XMM-Newton EPIC-pn spectrum is binned with a minimum of 15 photons. However, a single multicolor disk or blackbody model does not provide a satisfactory fit, leading to a large value. Next, we fit the spectrum using an absorbed power-law model, tbabs*zashift*powerlaw. The fit converges to an index (/dof=72.15/57), indicating a soft X-ray spectrum. However, this model did not account for the possible thermal emission from the disk. Therefore, we tried another model tbabs*zashift*(diskbb+powerlaw). This model produced a disk temperature of and a photon index of with /dof=64.61/55. Taking into account the potential intrinsic absorption, we tried to add a photoelectric absorbed component with tbabs*zashift*wabs*(diskbb+powerlaw) in the fitting. The parameters derived for this model were a disk temperature of , a photon index of , and a hydrogen density of (/dof=60.57/54). The power-law component could originate from nonthermal emission from the corona above the disk. To determine whether the absorbed component is necessary, we performed the F-test on the results and derived the F value of 3.6 and the P-value of 0.06. Therefore, we need to be careful about whether or not we need to include this absorbed component. The result is depicted in Figure 5. Although both a power-law model and diskbb model can fit the XRT spectrum well, the latter appears to be more plausible for TDEs because the soft X-ray emission in TDEs likely comes from the disk, which can be better characterized by diskbb. For the XMM spectrum, the additive power-law component may suggest the presence of a corona.
@@ 3.3 Host Galaxy Properties
We collected multiband photometry for the host galaxy of AT 2023lli from several archives, including Swift/UVOT, PanSTARRS, the Two Micron All Sky Survey, and Wide field Infrared Survey Explorer. We used the Python package Code Investigating GALaxy Emission (CIGALE; Boquien et al. 2019) to model the spectral energy distribution (SED) of the host galaxy. CIGALE can fit the SED of a galaxy from far-UV to radio and estimate its physical properties through the analysis of likelihood distribution. We assumed a delayed star formation history with an optional exponential burst and used the single stellar population of Bruzual & Charlot (2003). We adopted the dust attenuation curve of Calzetti et al. (2000). Additionally, dust emission is modeled using Dale et al. (2014), and active galactic nucleus (AGN) emission is calculated with the model of Stalevski et al. (2012, 2016). The fitting result is shown in Figure 6, which is well fitted solely by stellar components with zero contribution from the AGN component.
@@ 4.1.1 Comparison with Bumps in Other TDEs
Early bumps in the UV/optical light curves are not rare in TDEs that have a well-covered rising phase light curve (Wang et al., 2023b). In Figure 7, we compare the bump observed in AT 2023lli with previously reported cases for TDEs or TDE candidates. ASASSN-19bt, AT 2019azh, and AT 2020wey all display bumps that occur within 35 days of the main UV/optical peak (Holoien et al., 2019; Charalampopoulos et al., 2023; Faris et al., 2023). Wang et al. (2023b) reported early bumps that last around 10 days and occur within 30 days from the primary peak in ASASSN-18ap, AT 2019mha, and AT 2019qiz. The latter authors discussed several possibilities for the origin of the bumps, including the emission of unbound debris, the stream-stream collision, the vertical shock compression during the first passage, and the shock breakout of the debris collision. In contrast, the pTDE candidate ASASSN-14ko recently showed noticeable bumps before the main UV/optical peaks and rebrightenings after that. This can be interpreted as a consequence of the collision between the stream debris and the extended disk in the partial TDE scenario (Huang et al., 2023b), which distinguishes it from AT 2023lli and other TDEs.
@@ 4.1.2 Self-crossing of Stream Debris during Circularization and Delayed Accretion?
Thanks to the prominent optical early bump, AT 2023lli has been extensively observed in multiwavelength observations including Swift/XRT, making it a unique object with the earliest X-ray observations among all optical TDEs, starting from 40 days (rest-frame) prior to its peak. It is worth noting that ASASSN-19bt, the only TDE known to possess X-ray observations as early as 1 month before the peak, also exhibits a tentative early bump in its bolometric light curve (Holoien et al. 2019, see also Figure 7). Interestingly, neither TDE is X-ray detectable throughout the entire rising phase, yet AT 2023lli shows delayed X-ray emission in the late declining stage.
We propose a two-phase scenario to explain the optical and X-ray behaviors observed in this unique TDE. A chronological scheme is illustrated in Figure 8. A star is initially tidally disrupted into stream debris, which then undergoes a circularization process to form the accretion disk. Apsidal precession causes streams to self-cross, producing UV/optical emission (Piran et al., 2015). This process may explain the early bump observed in AT 2023lli. A new disk forms from the stream debris as it loses energy in the stream-stream collision and circularizes, resulting in the main UV/optical outburst. However, the X-rays from the accretion disk are blocked by the reprocessing layer, probably in the form of collision-induced outflow in the early stage (Lu & Bonnerot, 2020) and super-Eddington-accretion-induced outflow around peak (Metzger & Stone, 2016; Roth et al., 2016; Dai et al., 2018).
@@ 4.2 The Origin of the Episodic X-Ray Emission
Taking into account the Keplerian rotation of the optically thin region, its period is , where is the semi-major axis and is the gravitational constant. Assuming the time interval of the two initial X-ray emissions as the period, we can estimate the location of the optically thin region as
This result is of the same order of magnitude as the blackbody radius derived from UV/optical SED fitting, indicating that the “X-ray leak” is located near the outer boundary of the obscured materials. The initial appearance of the X-rays lasted only 2 days, the second lasted 11 days, and the third lasted much longer. This leads us to speculate that if the “X-ray leak” is from the same region, the area is expanding, and we can expect to observe longer-lasting X-ray emission in the future. Subsequent observations confirm this conjecture. A similar scenario was previously used to explain the variability of RXJ1301.9+2747 (Middleton & Ingram, 2015).
@@ 5 Conclusions
The advanced modern time-domain surveys are not only discovering TDEs with an increasing speed but also unveiling the diversity of TDEs in their temporal evolution. AT 2023lli, a nearby TDE at redshift 0.036, is exactly such a new member that exhibits notable features in its optical and X-ray light curves. In this study, we conducted continuous multiwavelength observations of it. These observations revealed characteristic properties, including an almost constant blackbody temperature, very broad Balmer and possible He II 4686 Å emission. However, it also exhibits some distinctive features that have captured our attention. The UV/optical light curves of AT 2023lli show a prominent bump that lasts nearly a month in the early rising phase. On the other hand, the time interval between the bump peak and the primary peak is nearly two months, making both timescales the longest among all TDEs with detected bumps, albeit only a handful so far. X-ray emission was not significantly detected until the late UV/optical decline stage ( days postpeak), despite two weak episodic emissions earlier. The XMM-Newton ToO observation of AT 2023lli, triggered by us upon persistent X-ray detection, revealed X-ray spectra with a possible intrinsic absorption component, albeit with large uncertainties. It is worth noting that AT 2023lli exhibits the power-law decay with an index of -4.10, making it one of the fastest-declining optically selected TDEs so far.
Taking all these observations into account, we suggest that the UV/optical bump may be caused by the self-crossing of the debris streams of the TDE, while the delayed accretion produces the UV/optical main outburst. The self-crossing or circularization of the debris streams produce outflows, which surround the accretion disk and cover it, leading to the absorption of soft X-ray photons from the inner disk and the detection of only the UV/optical emission. Due to the inhomogeneous distribution of the obscuring material, some regions were optically thin, allowing some X-ray photons to “leak” in the decline phase of the UV/optical. As the obscuring material evolved to be optically thin, the X-ray photons could be gradually detected. We have also discussed the possibility of double TDEs, in which the two stars in the binary are disrupted in sequence by the SMBH, which is also a possible scenario.
@@
We thank the anonymous referee for providing valuable comments, which helped to improve the manuscript. This work is supported by National Key Research and Development Program of China (2023YFA1608100), the National Natural Science Foundation of China (grants 12393814,12233008,12073025,12192221,12025303), the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0550200, XDB41000000, XDB0550300), Anhui Provincial Natural Science Foundation (2308085QA32), the China Manned Space Project (No. CMS-CSST-2021-A13) and the Fundamental Research Funds for Central Universities (WK3440000006). L.L. acknowledges the support of the National Natural Science Foundation of China (grant No. 12303050), the China Postdoctoral Science Foundation (2023M743397), and the Fundamental Research Funds for the Central Universities. The authors appreciate the support of the Cyrus Chun Ying Tang Foundations. The authors express their gratitude to Ping Cai for her help in creating the figure. We also appreciate the members of the WFST operation and maintenance team for their support. The Wide Field Survey Telescope (WFST) is a joint facility of the University of Science and Technology of China, Purple Mountain Observatory. This research uses data obtained through the Telescope Access Program (TAP), which has been funded by the TAP member institutes. Observations with the Hale Telescope at Palomar Observatory were obtained as part of an agreement between the National Astronomical Observatories, the Chinese Academy of Sciences, and the California Institute of Technology. The authors acknowledge the support of the Lijiang 2.4m telescope staff. Funding for the telescope has been provided by the Chinese Academy of Sciences and the People’s Government of Yunnan Province”. The authors acknowledge the use of public data from the Swift data archive. The authors thank the Swift ToO team for accepting our proposal and executing the observations. Based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. This work makes use of observations from the Las Cumbres Observatory global telescope network. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile.